OpenStudy (geekfromthefutur):

Which graph below shows a system of equations with infinitely many solutions? coordinate plane with two parallel lines coordinate plane with two intersecting lines coordinate plane with two coinciding lines coordinate plane with two intersecting lines

OpenStudy (cwrw238):

its the third option there are infinite number of points on the coinciding line

OpenStudy (geekfromthefutur):

oh ok thank you :3 can u help me with 2 more ?? ^-^

OpenStudy (cwrw238):

ok

OpenStudy (geekfromthefutur):

Sally has only nickels and dimes in her money box. She knows that she has less than $20 in the box. Let x represent the number of nickels in the box and y represent the number of dimes in the box. Which of the following statements best describes the steps to graph the solution to the inequality in x and y? Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 10x + 5y = 2000, and shade the portion above the line for positive values of x and y. Draw a dashed line to represent the graph of 5x + 10y = 2000, and shade the portion below the line for positive values of x and y. Draw a dashed line to represent the graph of 10x – 5y = 2000, and shade the portion below the line for positive values of x and y. A system of equations is shown below: 3x – y = 2 x + y = 6 The x-coordinate of the solution to this system of equations is _____. Numerical Answers Expected! Answer for Blank 1:

OpenStudy (cwrw238):

for the second question: add the 2 equations - this will eliminate y find the value of x then substitute this value of x into any one of the original equations to find y

OpenStudy (cwrw238):

as for the first one I dont know the value of a dime or a nickel

OpenStudy (geekfromthefutur):

oh thts cool :3 but would u at anychance know wat the the answer would be for for question 2??

OpenStudy (anonymous):

Which graph below shows a system of equations with infinitely many solutions? coordinate plane with two parallel lines coordinate plane with two intersecting lines coordinate plane with two coinciding lines coordinate plane with two intersecting lines

OpenStudy (secret-ninja):

Dime= 10 American cents Nickel= 5 American cents